# Towards Optimal Environmental Policies: Policy Learning under Arbitrary Bipartite Network Interference

### v2

@@ 2.3 Optimal Policy Learning using Parametric, Linear Exposure Models
In this section, we propose Q-Learning and A-Learning methods for solving Equation 4. Motivated by domain knowledge that outcome unit potential outcomes are determined by an aggregate exposure level, we assume the outcome model is linear in $\bar{A}_{i}$. For illustration purposes, we present methods based on parametric forms. Please note that such results can be generalized. Namely, the methods below can be extended to multiple timepoints, semi/nonparametric models under appropriate smoothness conditions, and general outcome model forms. See Supp. F for details. Here, we assume the following outcome model.
@@
for $\boldsymbol{\theta}_{0}=(\boldsymbol{\alpha}_{0},\boldsymbol{\beta}_{0})\in\Theta\subseteq\mathbb{R}^{|\boldsymbol{\theta}_{0}|}$, which parameterize $f_{0}(\mathbf{X}^{out}_{i};\boldsymbol{\alpha}_{0})$ and $f_{A}(\mathbf{X}^{out}_{i};\boldsymbol{\beta}_{0})$ respectively, and mean zero independent random variable $\epsilon_{i}$. Above, $f_{0}$ captures the main effect of the covariates and $f_{A}$ captures the treatment effect. By Equation 5, we can express the objective function as follows:
@@
Hence the effect of treating an intervention unit $j$ is fully captured by $\texttt{TotalEffect}_{j}$, precluding the need for defining a direct or spillover effect or making clustered interference assumptions. The quantity in Equation 6 will serve as a key object of interest throughout our analyses below.
Optimal Policy under a Linear Exposure Model.
@@
We are now set up to state our Q-Learning and A-Learning approach to obtaining statistical inference procedures. At a high-level, Q-Learning and A-Learning posit regression models for the mean outcome and determine the optimal policy by using a plug-in estimate for the decision rule. Q-Learning relies on a correctly-specified outcome regression model, while A-Learning is a variant of Q-Learning that fits a modified estimating equation to endow double robustness (Schulte et al., 2014). In our case of Equation 5, Q-Learning requires $\mu=(f_{0},f_{A})$ to be correctly specified to obtain the optimal policy and inference. As we will see in our approach below, A-Learning requires $f_{A}$ to be correctly specified but only requires correct specification of only one of baseline model $f_{0}$ or propensity score model $e$ to obtain valid inference and the optimal policy.
@@ 2.3.1 Q-Learning
Parametric Q-Learning proceeds by proposing a model $\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\bar{A}_{i};\boldsymbol{\theta}_{0}]$, and solving Equation 4 using the plug-in estimate $\hat{\boldsymbol{\theta}}$.
Estimation.
@@
Inference.
Under suitable regularity conditions, we find for covariance matrix $\Sigma$ defined in Section A.2,
[…]
@@ 2.3.2 A-Learning
Q-Learning requires the full model $\mu=(f_{0},f_{A})$ to be correctly specified in order to obtain the optimal policy and inference. In this section, we propose a doubly robust estimator for $\boldsymbol{\beta}$. Below, we are explicit about the dependence on $(\boldsymbol{\alpha},\boldsymbol{\beta})$ since they are estimated using different estimating equations. Let the propensity score for intervention unit $j$, $e_{j}$, be parameterized by $\boldsymbol{\gamma}_{0}\in\Gamma\subseteq\mathbb{R}^{|\boldsymbol{\gamma}_{0}|}$.
$e_{j}=\mathbb{P}(A_{j}=1\mid\mathbf{X}^{int}_{j};\boldsymbol{\gamma}_{0}).$ (9)
@@
Remark 2.1.
In Equation 10, the choice of $\lambda$ is flexible. For example, one can take this to be $\lambda(\mathbf{X}^{out}_{i},\mathbf{H}_{i};\boldsymbol{\beta})=c_{i}\frac{\partial}{\partial\boldsymbol{\beta}}f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta})$ for $c_{i}\in\mathbb{R}$. This is similar to (Leon et al., 2003; Robins, 2004; Schulte et al., 2014), in causal inference literature on defining a ‘class’ of estimating equations that may satisfy consistent estimation, asymptotic normality, and double-robustness. During our simulations and analysis, we take $c_{i}=\frac{1}{J}\sum_{j=1}^{J}H_{ij}$.
This formulation yields an approach to estimating $\boldsymbol{\beta}$ that is doubly-robust, as shown by the lemma below; if either $e$ or $f_{0}$ is correctly specified, we have consistent parameter estimation for $\beta$ (Bang & Robins, 2005). We clarify that the following approach still requires $f_{A}$ to be correctly specified.
@@
If $e$ or $f_{0}$ is correctly specified, then under the identification conditions (A1)-(A4),
@@
Theorem 1 (Asymptotic Normality).
Under suitable regularity conditions and (A11), that $\frac{J}{n}\rightarrow\mathcal{R}$ a fixed constant, as $J,n\rightarrow\infty$, we have for covariance matrix $\Omega$ defined in Section A.3.1,
[…]
@@ 2.4 Estimating the Optimal Policy
With the treatment effect estimates in hand via $\hat{\boldsymbol{\beta}}$ from either Q-Learning or A-Learning, we now describe how to estimate the optimal policy.
Unconstrained Optimal Policy.
From Equation 7, we use the plug-in $\hat{\boldsymbol{\beta}}$ to find:
$\hat{\pi}_{j}^{*}(\widehat{\texttt{TotalEffect}}(\hat{\boldsymbol{\beta}}))=\textbf{1}\{\widehat{\texttt{TotalEffect}}(\hat{\boldsymbol{\beta}})<0\}$ (13)
Cost-Constrained Optimal Policy.
The constrained optimal policy is given by solving Equation 4 with plug-in $\hat{\boldsymbol{\beta}}$. If we assume that the costs and cost constraints are positive (i.e. $c_{j},C>0$), we can obtain an efficient $O(J)$ computable solution to Equation 4 via the fractional knapsack solution, taking $c_{j}$ as weights and $-Y_{i}$ as profits (see Ch. 15 Cormen et al. (2009)). To translate this to our setup, we define the following:
@@
This quantity represents the benefit-cost ratio of installing a scrubber at a power plant $j$ relative to the cost. Such “relative benefits” play a key role in determining cost-constrained optimal intervention strategies, as $\texttt{TotalEffect}_{j}$ does for the unconstrained counterpart. We will estimate this ratio using our model output to explain and contextualize our findings in Section 4.4.2.
@@ 3 Simulation Study
In this section, we conduct an empirical Monte Carlo study to validate the methods in Section 2. The covariates and interference matrix are the same as from our data application across all simulations. Next to these “empirical variables”, we randomly generate treatment assignments and outcomes according to a fixed Signal-to-Noise Ratio (SNR). The Bias of $\boldsymbol{\beta}$ (Equation C.1), the RMSE of $\widehat{\texttt{TotalEffect}}$ (Equation C.2), and the coverage probabilities are reported across all simulations. The equations for the metrics and parameters used for simulation are found in Supp. C. Throughout simulations, we assume the following form for the true outcome and propensity score models:
@@
$\mathrm{log}\left(\frac{e_{j}}{1-e_{j}}\right)$ $=\gamma_{\mathrm{intercept}}+\mathbf{X}^{int}_{j}\boldsymbol{\gamma}_{1}+(\mathbf{X}^{int}_{j})^{2}\boldsymbol{\gamma}_{2}$ (17)
In the simulations, we consider scenarios with fully correct model specification and with $f_{0}$ and/or $e$ misspecified to demonstrate the double robustness property of A-learning. Misspecification of each component is induced by fitting models imposing linear relationships with covariates in a given component instead of quadratic. Specifically, outcome model misspecification is carried out by setting $\boldsymbol{\alpha}_{2}=0$. The misspecification of the propensity score model is introduced by setting $\boldsymbol{\gamma}_{2}=0$. Throughout, to construct an empirically grounded simulation framework that accurately reflects the real-world dynamics of our data application, $\boldsymbol{\gamma}$ is chosen so that simulated average of propensities $\frac{1}{J}\sum_{j=1}^{J}e_{j}$ is within 0.01 of the observed empirical average of treatments in the dataset (see final row of Table B.1). Further, $\boldsymbol{\theta}=(\boldsymbol{\alpha},\boldsymbol{\beta})$ is chosen so that the simulated average $\frac{1}{n}\sum_{i=1}^{n}Y_{i}$, based on the expected exposure level $\bar{A}=\frac{1}{J}\sum_{j=1}^{J}H_{ij}e_{j}$ from Step 2, is within 0.001 of the empirical average in the dataset (see penultimate row of Table B.2).
Our simulation setup is as follows.
@@
    Generate intervention unit treatments: $A_{j}\sim\mathrm{Bernoulli}(e_{j})$, where $e_{j}$ follows Equation 17.
@@
    where $\mathbb{V}(\epsilon_{i})=\mathbb{V}\left(\frac{\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\mathbf{H}_{i},\bar{A}_{i};\boldsymbol{\theta}]}{3}\right)$ and $f_{0}$, $f_{A}$ follow Equation 15 and 16 respectively.
@@
    where the Q-learning models are estimated using least squares regression and the A-learning models solve Equation 11, 12 and 10 jointly.
@@
Simulation Results
Method BS ($f_{0}$) PS ($e$) Bias RMSE Coverage
Q-Learning - 0.021 0.73 94.84
Q-Learning - 0.22 7.10 15.40
A-Learning 0.024 1.32 98.57
A-Learning 0.024 1.32 98.37
A-Learning 0.057 1.74 95.23
A-Learning 0.058 1.75 94.56
Table 1: Simulation Results over 1000 iterations. BS denotes baseline model, $f_{0}$, and PS denotes propensity score model, $\pi$. See the definitions for Bias (Equation C.1) and RMSE (Equation C.2) above.
The results from 1000 simulations are shown in Table 1. The correctly specified Q-Learning model achieves the lowest bias and RMSE, consistent with common knowledge that a correctly-specified outcome model tends to outperform other methods. A-Learning with correctly specified models is competitive with Q-Learning correctly specified. Under misspecification of the propensity score model ($e$), A-Learning gives nearly equivalent performance to A-Learning under a fully correct model specification (first and second rows of A-Learning). As we introduce the misspecification of the outcome model in Q-Learning and A-Learning, the error metrics increase.
Based on these simulations, under equivalent misspecification forms, the methodology seems less sensitive to propensity score model misspecification than outcome model misspecification. This phenomenon was also observed in simulations on doubly robust causal effect estimation under BNI in Chen, Bargagli-Stoffi, Kim & Nethery (2024). Regardless, under equivalent baseline model misspecification (second row of Q-Learning and third row of A-Learning), A-Learning results in lower estimation error than Q-Learning. A-Learning yields good empirical coverage for significance level $\alpha=0.05$, even with misspecified models. In contrast, under equivalent outcome model misspecification, Q-Learning has extremely low coverage, affecting inference more severely than A-Learning.
In general, we find a lower estimation error in treatment effect estimates and a higher coverage under various levels of misspecification in A-Learning compared to misspecified Q-Learning, validating the theoretical results shown above in Section 2.3.1 and Section 2.3.2.

### v3

@@ 2.3 Optimal Policy Learning using Parametric, Linear Exposure Models
In this section, we propose Q-Learning and A-Learning methods for solving Eq. 4. Motivated by domain knowledge that outcome unit potential outcomes are determined by an aggregate exposure level, we assume the outcome model is linear in $\bar{A}_{i}$. This assumption not only simplifies the policy optimization, but enhances the interpretability of the optimal solution. For illustration purposes, we present methods based on parametric forms. Please note that such results can be generalized. Namely, the methods below can be extended to semi/nonparametric models under appropriate smoothness conditions and general outcome model forms. See Appendix Appendix P for details. At a high level, we are able to derive inferential results by employing the multiplier central limit theorem as Su et al. (2019), and regret bounds by spatial mixing assumptions of the outcome units as Jenish and Prucha (2009). We now describe our setup in more detail, while the formal proofs can be found in the appendix. Here, we assume the following outcome model.
@@
for $\boldsymbol{\theta}_{0}=(\boldsymbol{\alpha}_{0},\boldsymbol{\beta}_{0})\in\Theta\subseteq\mathbb{R}^{|\boldsymbol{\theta}_{0}|}$, which parameterize $f_{0}(\mathbf{X}^{out}_{i};\boldsymbol{\alpha}_{0})$ and $f_{A}(\mathbf{X}^{out}_{i};\boldsymbol{\beta}_{0})$ respectively, and mean zero independent random variable $\epsilon_{i}$. Above, $f_{0}$ captures the main effect of the covariates and $f_{A}$ captures the treatment effect. By Eq. 5, we can express the objective function as follows:
@@
Hence the effect of treating an intervention unit $j$ is fully captured by $\texttt{TotalEffect}_{j}$, precluding the need for defining a direct or spillover effect or making clustered interference assumptions. The quantity in Eq. 6 will serve as a key object of interest throughout our analyses below.
2.3.1 Optimal Policy under a Linear Exposure Model.
@@
We are now set up to state our Q-Learning and A-Learning approach to obtaining statistical inference procedures. At a high-level, Q-Learning and A-Learning posit regression models for the mean outcome and determine the optimal policy by using a plug-in estimate for the decision rule. Q-Learning relies on a correctly-specified outcome regression model, while A-Learning is a variant of Q-Learning that fits a modified estimating equation to endow double robustness (Schulte et al., 2014). In our case of Eq. 5, Q-Learning requires $\mu=(f_{0},f_{A})$ to be correctly specified to obtain the optimal policy and inference. As we will see in our approach below, A-Learning requires $f_{A}$ to be correctly specified but only requires correct specification of only one of baseline model $f_{0}$ or propensity score model $e$ to obtain valid inference and the optimal policy.
@@ 2.3.2 Q-Learning
Parametric Q-Learning proceeds by proposing a model $\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\bar{A}_{i};\boldsymbol{\theta}_{0}]$, and solving Eq. 4 using the plug-in estimate $\hat{\boldsymbol{\theta}}$.
Estimation
@@
Inference
Under suitable regularity conditions, we find for covariance matrix $\Sigma$ defined in Appendix Appendix J.2,
$\sqrt{n}(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}_{0})\xrightarrow{d}N(0,\Sigma)$
The proof follows in a similar manner to Theorem 3 below invoking the multiplier central limit theorem Kosorok (2008). We defer the details to Appendix Appendix J.2.
@@ 2.3.3 A-Learning
Q-Learning requires the full model $\mu=(f_{0},f_{A})$ to be correctly specified in order to obtain the optimal policy and inference. In this section, we propose a doubly robust estimator for $\boldsymbol{\beta}$. Below, we are explicit about the dependence on $(\boldsymbol{\alpha},\boldsymbol{\beta})$ since they are estimated using different estimating equations. Let the propensity score for intervention unit $j$, $e_{j}$, be parameterized by $\boldsymbol{\gamma}_{0}\in\Gamma\subseteq\mathbb{R}^{|\boldsymbol{\gamma}_{0}|}$, or $e_{j}=\mathbb{P}(A_{j}=1\mid\mathbf{X}^{int}_{j};\boldsymbol{\gamma}_{0})$.
@@
Remark 3.
In Eq. 9, the choice of $\lambda$ is flexible. For example, one can take this to be $\lambda(\mathbf{X}^{out}_{i},\mathbf{H}_{i};\boldsymbol{\beta})=c_{i}\frac{\partial}{\partial\boldsymbol{\beta}}f_{A}(\mathbf{X}^{out}_{i},\boldsymbol{\beta})$ for $c_{i}\in\mathbb{R}$. This is similar to (Leon et al., 2003; Robins, 2004; Schulte et al., 2014), in causal inference literature on defining a ‘class’ of estimating equations that may satisfy consistent estimation, asymptotic normality, and double-robustness. During our simulations and analysis, we take $c_{i}=\frac{1}{J}\sum_{j=1}^{J}H_{ij}$.
Remark 4.
In our specification of the propensity score, we may augment $\mathbf{X}_{j}^{int}$ to incorporate information from $\mathbf{X}^{out}_{i},\mathbf{H}_{i}$ by including weighted summaries of outcome unit covariates, weighted by the interference level $\mathbf{H}_{i}$ (e.g., $\mathbf{X}^{out}_{j}=\frac{1}{n}\sum_{i=1}^{n}H_{ij}\mathbf{X}^{out}_{i}$). Such an approach is natural for incorporating outcome level covariates in the intervention model, up- or down-weighting the influence of outcome unit characteristics based on how much a given outcome unit is affected by the intervention unit in question.
This formulation yields an approach to estimating $\boldsymbol{\beta}$ that is doubly-robust, as shown by the lemma below; if either $e$ or $f_{0}$ is correctly specified, we have consistent parameter estimation for $\boldsymbol{\beta}$ (Bang and Robins, 2005). We clarify that the following approach still requires $f_{A}$ to be correctly specified, or the true treatment effects will not be recovered.
@@
If $e$ or $f_{0}$ is correctly specified, then under the identification conditions (I1)-(I4),
@@
Theorem 3 (Asymptotic Normality for A-Learning).
Under suitable regularity conditions and the assumption that $\frac{J}{n}\rightarrow\mathcal{R}$ a fixed constant, as $J,n\rightarrow\infty$, we have for covariance matrix $\Omega$ defined in Section Appendix J.4,
$\sqrt{n}\begin{bmatrix}\hat{\boldsymbol{\alpha}}-\boldsymbol{\alpha}_{0}\\ \hat{\boldsymbol{\beta}}-\boldsymbol{\beta}_{0}\end{bmatrix}\xrightarrow{d}N(0,\Omega)$
@@ 2.4 Estimating the Optimal Policy
With the treatment effect estimates in hand via $\hat{\boldsymbol{\beta}}$ from either Q-Learning or A-Learning, we now describe how to estimate the optimal policy. In Appendix Appendix O, we describe how to utilize our learned treatment effect parameters to perform policy learning on future, possibly independent samples.
2.4.1 Unconstrained Optimal Policy.
From Eq. 7, we use the plug-in $\hat{\boldsymbol{\beta}}$ to find:
$\hat{\pi}_{j}^{*}(\texttt{TotalEffect}(\hat{\boldsymbol{\beta}})_{j})=\textbf{1}\{\texttt{TotalEffect}(\hat{\boldsymbol{\beta}})_{j}<0\}$ (12)
2.4.2 Cost-Constrained Optimal Policy.
The constrained optimal policy is given by solving Eq. 4 with plug-in $\hat{\boldsymbol{\beta}}$. If we assume that the costs and cost constraints are positive (i.e. $c_{j},C>0$), we can obtain an efficient $O(J)$ computable solution to Eq. 4 via the fractional knapsack solution, taking $c_{j}$ as weights and $Y_{i}$ as profits (see Ch. 15 Cormen et al. (2009)). To translate this to our setup, we define the benefit-cost ratio of installing a scrubber at a power plant $j$ relative to the cost, as
@@
Such “relative benefits” play a key role in determining cost-constrained optimal intervention strategies, as $\texttt{TotalEffect}_{j}$ does for the unconstrained counterpart. We will estimate this ratio using our model output to explain and contextualize our findings in Section 4.4.2.
2.5 Regret Bound Analysis
In the context of policy learning, it is key to study the regret of the estimated optimal policy with respect to the true optimal policy. These regret bounds guarantee that our policy learned acts optimally as our sample size grows, and characterizes the rate at which optimality is achieved.
Since the optimal policy relies on estimating $\texttt{TotalEffect}_{j}$, we first show consistency and estimation rates for this quantity. With this in hand, regret bounds follow. For brevity, let $\hat{\eta}_{j}=\texttt{TotalEffect}(\hat{\boldsymbol{\beta}})_{j}$ and $\eta_{j}=\texttt{TotalEffect}(\boldsymbol{\beta})_{j}$. Throughout, we rely on a spatial setup and mixing assumptions that are standard in spatial statistics literature, similar to that in Jenish and Prucha (2009). The full setup and assumptions are described in Section Appendix J.5. To provide some more context into our setup, (S1) formalizes how we observe spatial units, saying that asymptotically, the number of units we observe grows in space, while (S2) says that units that are distance $r$ apart have dependency that decays sufficiently fast.
These assumptions are well-suited for our scientific setting, where interference from air pollution exposure exhibits spatial spillovers that are not adequately captured by traditional finite-clustering assumptions. With these in hand, we find the following.
Theorem 4 (Consistency of Treatment Effect Estimation).
Assume spatial mixing (S1)-(S2), and suitable regularity conditions. Then we can estimate the treatment effect function for intervention $j$ at $\sqrt{n}$ rates, or
$\mathbb{E}\frac{1}{n}|\hat{\eta}_{j}-\eta_{j}|\lesssim\sqrt{\frac{1}{n}}$
With this in hand, we may conclude that our regret is asymptotically 0. I.e.
Theorem 5 (Convergence to the Optimal Policy).
Assume spatial mixing (S1)-(S2), and suitable regularity conditions. Denote our estimated policy from utilizing $\hat{\eta}_{j}$ by $\hat{\pi}^{*}$. Then our policy is asymptotically optimal (with regret approaching 0), or
$|\mathbb{E}_{\hat{\pi}^{*}}[Y_{i}(\bar{a}_{i})]-\mathbb{E}_{\pi^{*}}[Y_{i}(\bar{a}_{i})]|\xrightarrow{p}0$
@@ 3 Simulation Study
In this section, we conduct an empirical Monte Carlo study to validate the methods in Section 2. The covariates and interference matrix are the same as from our data application across all simulations. Next to these “empirical variables”, we randomly generate treatment assignments and outcomes according to a fixed Signal-to-Noise Ratio (SNR). The Bias of $\boldsymbol{\beta}$ (Eq. Appendix L.1), the RMSE of $\widehat{\texttt{TotalEffect}}$ (Eq. Appendix L.2), and the coverage probabilities are reported across all simulations. The equations for the metrics used in the simulation are found in Appendix Appendix L. Throughout simulations, we assume the following form for the true outcome and propensity score models:
@@
$\log\left(\frac{e_{j}}{1-e_{j}}\right)$ $=\gamma_{\mathrm{intercept}}+\mathbf{X}^{int}_{j}\boldsymbol{\gamma}_{1}+(\mathbf{X}^{int}_{j})^{2}\boldsymbol{\gamma}_{2}$ (16)
In the simulations, we consider scenarios with fully correct model specification and with $f_{0}$ and/or $e$ misspecified to demonstrate the double robustness property of A-learning. Misspecification of each component is induced by fitting models imposing linear relationships with covariates in a given component instead of quadratic. Specifically, outcome model misspecification is carried out by setting $\hat{\boldsymbol{\alpha}}_{2}=0$. The misspecification of the propensity score model is introduced by setting $\hat{\boldsymbol{\gamma}}_{2}=0$. Throughout, to construct an empirically grounded simulation framework that accurately reflects the real-world dynamics of our data application, $\boldsymbol{\gamma}$ is chosen so that simulated average of propensities $\frac{1}{J}\sum_{j=1}^{J}e_{j}$ is within 0.01 of the observed empirical average of treatments in the dataset (see final row of Table Appendix K.1). Further, $\boldsymbol{\theta}=(\boldsymbol{\alpha},\boldsymbol{\beta})$ is chosen so that the simulated average $\frac{1}{n}\sum_{i=1}^{n}Y_{i}$, based on the expected exposure level $\bar{A}=\frac{1}{J}\sum_{j=1}^{J}H_{ij}e_{j}$ from Step 2, is within 0.01 of the empirical average in the dataset (see penultimate row of Table Appendix K.2).
Our simulation setup is as follows:
@@
    Generate intervention unit treatments: $A_{j}\sim\mathrm{Bernoulli}(e_{j})$, where $e_{j}$ follows Eq. 16.
@@
    where $\mathbb{V}(\epsilon_{i})=\mathbb{V}\left(\frac{\mathbb{E}[Y_{i}\mid\mathbf{X}^{out}_{i},\mathbf{H}_{i},\bar{A}_{i};\boldsymbol{\theta}]}{3}\right)$ and $f_{0}$, $f_{A}$ follow Eqs. 14 and 15 respectively.
@@
    where the Q-learning models are estimated using least squares regression and the A-learning models solve Eqs. 10, 11 and 9 jointly.
@@
The results from 1000 simulations are shown in Table 1. The correctly specified Q-Learning model achieves the lowest estimation error, consistent with common knowledge that a correctly-specified outcome model tends to outperform other methods. A-Learning with correctly specified models is competitive with Q-Learning correctly specified. Under misspecification of the propensity score model ($e$), A-Learning gives nearly equivalent performance to A-Learning under a fully correct model specification (first and second rows of A-Learning). As we introduce misspecification of the outcome model in Q-Learning and A-Learning, the error metrics increase.
Based on these simulations, under equivalent misspecification forms, the methodology seems less sensitive to propensity score model misspecification than outcome model misspecification. This phenomenon was also observed in simulations on doubly robust causal effect estimation under BNI in Chen et al. (2024b). Regardless, under equivalent baseline model misspecification (second row of Q-Learning and third row of A-Learning), A-Learning results in lower estimation error than Q-Learning. A-Learning yields good empirical coverage for significance level $\alpha=0.05$, even with misspecified models. In contrast, under equivalent outcome model misspecification, Q-Learning has extremely low coverage, affecting inference more severely than A-Learning.
In general, we find a lower estimation error and a higher coverage under various levels of misspecification in A-Learning compared to misspecified Q-Learning, validating the theoretical results shown above in Section 2.3.2 and Section 2.3.3.